OpenStudy (anonymous):

what does the square root of p^6 simplify down to?

OpenStudy (anonymous):

\[\sqrt{p}=p^{1/2}\]; \[\sqrt{p^6}=p^{6/2}=p^3\]

OpenStudy (anonymous):

thank you